OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 3x + 1 A. (6, 18) and (0, 1) B. (19, 6) and (0, 1) C. (6, 19) and (0, 1) D. (6, 19) and (0, 0)

OpenStudy (anonymous):

Do you know how this works? You replace the x with the first number of the pair, and replace y with the second number of the pair so if the pair was (1,2), then you would replace x with 1, and y with 2. Make sense?

OpenStudy (anonymous):

Welcome to Open Study!

OpenStudy (anonymous):

is it b?

OpenStudy (anonymous):

Sorry I meant 6 = 3(19) + 1

OpenStudy (anonymous):

o its C!

OpenStudy (anonymous):

@tutuyy Is this equation true? 6 = 3(19) + 1?

OpenStudy (anonymous):

Yep! Great Job!

OpenStudy (anonymous):

YESA!

OpenStudy (anonymous):

so it is c right? @HailKK

OpenStudy (anonymous):

Yep! I already tested it out like 5 minutes ago.

OpenStudy (anonymous):

Which equation is shown on the graph? A. y = 2x B. y = 2x + 2 C. y = x + 2 D. y = x – 2

OpenStudy (anonymous):

@HailKK

OpenStudy (anonymous):

Give me a second

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Ugh I hate these ones...

OpenStudy (anonymous):

These things are hard for me to explain but I'll try...

OpenStudy (anonymous):

ill fan you

OpenStudy (anonymous):

Does any point in the line on the graph represent that y is 2 times the value of x? If so, A is the answer. If not cross out A

OpenStudy (anonymous):

whats the answer? A?

OpenStudy (anonymous):

No. Because it does not apply to any point on the line. Only one.

OpenStudy (anonymous):

Whats the answer?

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yes buy whays the answer?

OpenStudy (anonymous):

Look at point 1,3 of the line on the graph. What does this show about the relation between x and y?

OpenStudy (anonymous):

y is always higher than x by how much?

OpenStudy (anonymous):

ITS B!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Does 1(2) +2 = 3? Because that is what you are implying

OpenStudy (anonymous):

ITS C!@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@@

OpenStudy (anonymous):

Yep! because on any point of the graph, x+2 = y

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

@tutuyy